# Gravitational mode mixing around black holes in scalar-tensor theories with parity-violating terms

### v1

@@ V Rotating black hole solutions without $\mathbb{Z}_{2}$ symmetry
where the background solutions $\bar{\phi}$ and $\bar{g}_{\mu\nu}$ are given by Eqs. (4)-(5) with Eqs. (7)-(14), and the form of $\delta g_{\mu\nu}$ is given by
@@
Solving the equations of motion (2) and (3) with the ansatz (62)–(67) perturbatively, we obtain slowly rotating black hole solutions. The lowest order solution ${\cal O}(\epsilon^{0})$ corresponds to the slowly rotating Kerr black hole up to the first order of the spin parameter $\chi$
@@
Note again that ${\cal O}(\epsilon^{0})$ term in Eq. (80) represents the the slowly rotating Kerr solution. These slowly rotating solutions violate the $\mathbb{Z}_{2}$ symmetry $\theta\to\pi-\theta$ through the $\cos{\theta}$ terms in (78) and (79). This kind of $\mathbb{Z}_{2}$ violation can be seen in the purely tensorial gravity theories with higher curvature corrections Cardoso et al. (2018); Cano and Ruipérez (2019) and parity-violating modified gravity Cunha et al. (2018); Tahara et al. (2023).
@@ VI Mode mixing in tidal response
where $C^{\rm s}$ and $C^{\pm}$ are integration constants, corresponding to the strength of the external tidal fields. The absence of decaying modes in the far region in Eqs. (81)-(83) implies vanishing Love numbers around Schwarzschild black holes Binnington and Poisson (2009). To determine the non-GR effect, we expand the static solutions of Eq. (27) as
@@
where $\zeta(3)$ is the Riemann zeta function $\zeta(n)$ with $n=3$. This result shows that an even parity external tidal field $C^{+}$ induces an odd parity gravitational response in $\tilde{\Psi}^{-}_{\rm 2nd}$, and vice versa for the odd parity external tidal field $C^{-}$, as seen in Cardoso et al. (2018). However, due to the presence of the logarithmic terms in Eqs (85) and (86), it is not obvious how to define the tidal Love number.111111 Note that the logarithmic terms appear in the world-line EFT of a black hole Kol and Smolkin (2012); Hui et al. (2021); Ivanov and Zhou (2023); Charalambous et al. (2022); Saketh et al. (2023); Mandal et al. (2024) even in GR, modified gravity theories Creci et al. (2023), and gravitational EFTs Charalambous et al. (2022); De Luca et al. (2023); Katagiri et al. (2024). In the world-line EFT, it is argued that these logarithmic terms can be cancelled out by taking into account ultraviolet physics in concrete models, and the tidal Love number becomes well-defined Ivanov and Zhou (2023). Nevertheless, the tidal response through the even-odd gravitational couplings can be suggested. The origin of the number $\zeta(3)$ in Eqs (85) and (86) is the regularity condition at the horizon, so the asymptotic forms of $\tilde{\Psi}^{-}_{\rm 2nd}$ and $\tilde{\Psi}^{+}_{\rm 2nd}$ probably contain information about the horizon. This suggests that $\tilde{\Psi}^{-}_{\rm 2nd}$ and $\tilde{\Psi}^{+}_{\rm 2nd}$ contain decaying modes, i.e., the non-vanishing tidal responses, because purely growing modes, which are determined only from the behaviour around the asymptotic region $r\to\infty$, are not expected to contain information about the horizon.
@@ VII Summary and discussion
Finally, we discuss the generality of even-odd gravitational couplings in scalar-tensor theories. If we assume that the scalar field $\phi$ has a background value $\bar{\phi}$, the dCS gravity sector in the Lagrangian behaves
@@
where $h^{\pm}$ are the odd and even-type gravitational perturbations, respectively. This demonstrates the existence of the even-odd gravitational couplings when $\bar{\phi}\neq 0$. This structure is common in scalar-tensor theories with a single scalar field coupled to a parity-violating term other than the Pontryagin term. While the even-odd gravitational couplings have already appeared in higher curvature gravity theories Cardoso et al. (2018); de Rham et al. (2020); Cano et al. (2022b, 2023a), such couplings generically arise in scalar-tensor theories with parity-violating terms. The analysis of even-odd gravitational couplings in the general setting is left for future work Shin’ichi et al. .
@@ Acknowledgements.
We are grateful for fruitful discussions with Takuya Katagiri, and Kazufumi Takahashi. This work was supported by the research fellowship at the National Institute of Technology, Oyama College (S.H.), JSPS KAKENHI Grant Nos. JP22K03626 (M.K.) and JP21H01080 (M.Y.), and IBS under the project code IBS-R018-D3 (M.Y.). We acknowledge the hospitality at APCTP where part of this work was done during the focus research program “Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”.
@@ Appendix A Explicit forms of master variables, deviations of sound speed and potentials
In Eqs. (27)–(46), the coupled master equations of the scalar, odd, and even gravitational modes are derived. Due to the couplings, the master variables are mixed with other modes at higher orders in $\epsilon$. The relations of the master variables in the coupled master equation (27) and those in Eqs. (20)-(22) are given by
@@
$-\frac{8ir_{\rm g}^{4}\omega(75r^{2}+78rr_{\rm g}-10r_{\rm g}^{2})}{15r^{5}}\frac{{\rm d}\tilde{\Psi}^{-}}{{\rm d}r}\Bigg{\}}$ (104)
@@ Appendix B Perturbative eigenvalues of symmetric matrix
In this section, we study the eigenvalues of symmetric matrices that are diagonal at the lowest order but have non-diagonal components perturbatively. Since the characteristics of the perturbative spectrum of eigenvalues of symmetric matrices are similar to those of the spectrum of quasi-normal frequencies of black holes, this study is useful for understanding the results of Sec. IV.
@@ B.2 $2\times 2$ case: degenerate
Substituting $\lambda=2+\Delta\lambda$ into Eq. (118), we obtain $(\Delta\lambda)^{2}-\delta^{2}=0$. This implies that the eigenvalues are $\lambda=2\pm\delta$. For symmetric matrices that are degenerate at the lowest order, the correction of the eigenvalue appears at ${\cal O}(\delta)$. This is because the leading correction term of $\Delta\lambda$ in the characteristic equation is ${\cal O}((\Delta\lambda)^{2})$ if the lowest order spectra are degenerate.
@@ B.3 $3\times 3$ case
If $p$ takes different values when $b_{1}=0$ and when $b_{2}=0$ while keeping the condition $b_{1}b_{2}+b_{3}=0$, the branches of Eq. (131) and Eq. (132) exchange for $b_{1}=0$ and $b_{2}=0$ cases. This phenomenon can be seen in the QNM spectra for our setup.
@@ Appendix C Static solutions as tidal response for $\ell=2$
In Sec. VI, we discussed the tidal response from the new even-odd couplings. In this section, we present explicit forms of the static solutions with the regularity condition at the horizon for $\ell=2,\omega=0$ in the master equations (27). To obtain perturbative solutions in $\epsilon$, we expand the solutions around the lowest order solutions up to ${\cal O}(\epsilon^{2})$ as Eq. (84). The lowest-order solutions $\tilde{\Psi}^{\rm s}_{\rm Sch},~{}\tilde{\Psi}^{-}_{\rm Sch}$, and $\tilde{\Psi}^{+}_{\rm Sch}$ are given by Eqs. (81)-(83), which corresponds to the Schwarzschild case. We derive the correction terms by solving Eq. (27) with the ansatz in Eq. (84).
C.1 scalar field led case
@@
Through the scalar-odd coupling $V^{\rm s-}_{\rm 1st}$ and scala-even coupling $V^{\rm s+}_{\rm 1st}$ and the self-scalar coupling $V^{\rm s}_{\rm 2nd}$ in Eq. (46), the higher-order solutions are given by
@@
Let us consider solutions led by the odd parity external tidal field. The lowest-order growing odd solution is given by Eq (82)
@@
We study solutions led by the even parity external tidal field. The lowest-order growing solution is given by Eq (83)

### v2

@@ V Rotating black hole solutions without $\mathbb{Z}_{2}$ symmetry
where the background solutions $\bar{\phi}$ and $\bar{g}_{\mu\nu}$ are given by Eqs. (4) and (5) with Eqs. (7)–(14), and the form of $\delta g_{\mu\nu}$ is given by
@@
Solving the equations of motion (2) and (3) with the ansatz (47)–(52) perturbatively, we obtain slowly rotating black hole solutions. The lowest-order solution ${\cal O}(\epsilon^{0})$ corresponds to the slowly rotating Kerr black hole up to the first order of the spin parameter $\chi$
@@
Note again that ${\cal O}(\epsilon^{0})$ term in Eq. (65) represents the the slowly rotating Kerr solution. These slowly rotating solutions violate the $\mathbb{Z}_{2}$ symmetry $\theta\to\pi-\theta$ through the $\cos{\theta}$ terms in (63) and (64). This kind of $\mathbb{Z}_{2}$ violation can be seen in the purely tensorial gravity theories with higher-curvature corrections Cardoso et al. (2018); Cano and Ruipérez (2019) and parity-violating modified gravity Cunha et al. (2018); Tahara et al. (2023).
@@ VI Mode mixing in tidal response
where $C^{\rm s}$ and $C^{\pm}$ are integration constants, corresponding to the strength of the external tidal fields. The absence of decaying modes in the far region in Eqs. (66)–(68) implies vanishing Love numbers around Schwarzschild black holes Binnington and Poisson (2009). To determine the non-GR effect, we expand the static solutions of Eq. (27) as
@@
where $\zeta(3)$ is the Riemann zeta function $\zeta(n)$ with $n=3$. This result shows that an even parity external tidal field $C^{+}$ induces an odd parity gravitational response in $\tilde{\Psi}^{-}_{\rm 2nd}$, and vice versa for the odd parity external tidal field $C^{-}$, as seen in Cardoso et al. (2018). However, due to the presence of the logarithmic terms in Eqs (70) and (71), it is not obvious how to define the tidal Love number.131313 Note that the logarithmic terms appear in the worldline EFT of a black hole Kol and Smolkin (2012); Hui et al. (2021); Ivanov and Zhou (2023); Charalambous et al. (2022); Saketh et al. (2023); Mandal et al. (2024) even in GR, modified gravity theories Creci et al. (2023), and gravitational EFTs Charalambous et al. (2022); De Luca et al. (2023); Katagiri et al. (2024); Bhattacharyya et al. (2024). In the world-line EFT, it is argued that these logarithmic terms can be cancelled out by taking into account ultraviolet physics in concrete models, and the tidal Love number becomes well-defined Ivanov and Zhou (2023). Nevertheless, the tidal response through the even-odd gravitational couplings can be suggested. The origin of the number $\zeta(3)$ in Eqs (70) and (71) is the regularity condition at the horizon, so the asymptotic forms of $\tilde{\Psi}^{-}_{\rm 2nd}$ and $\tilde{\Psi}^{+}_{\rm 2nd}$ probably contain information about the horizon. This suggests that $\tilde{\Psi}^{-}_{\rm 2nd}$ and $\tilde{\Psi}^{+}_{\rm 2nd}$ contain decaying modes, i.e., the non-vanishing tidal responses, because purely growing modes, which are determined only from the behaviour around the asymptotic region $r\to\infty$, are not expected to contain information about the horizon.
@@ VII Summary and discussion
Finally, we discuss the generality of even-odd gravitational couplings in scalar-tensor theories. If we assume that the scalar field $\phi$ has a background value $\bar{\phi}$, the dCS gravity sector in the Lagrangian behaves as
@@
where $h^{\pm}$ are the odd and even-type gravitational perturbations, respectively. This demonstrates the existence of the even-odd gravitational couplings when $\bar{\phi}\neq 0$. This structure is common in scalar-tensor theories with a single scalar field coupled to a parity-violating term other than the Pontryagin term. While the even-odd gravitational couplings have already appeared in higher-curvature gravity theories Cardoso et al. (2018); de Rham et al. (2020); Cano et al. (2022b, 2023a), such couplings generically arise in scalar-tensor theories with parity-violating terms. The analysis of even-odd gravitational couplings in the general setting is left for future work Shin’ichi et al. .
@@ Acknowledgements.
We are grateful for fruitful discussions with Takuya Katagiri, and Kazufumi Takahashi. This work was supported by the research fellowship at the National Institute of Technology, Oyama College (S.H.), JSPS KAKENHI Grant No. JP22K03626 (M.K.), and No. JP21H01080 (M.Y.), and IBS under the Project No. IBS-R018-D3 (M.Y.). We acknowledge the hospitality at APCTP where part of this work was done during the focus research program “Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”.
@@ Appendix A Explicit forms of master variables, deviations of sound speed and potentials
In Eqs. (27)–(31), the coupled master equations of the scalar, odd, and even gravitational modes are derived. Due to the couplings, the master variables are mixed with other modes at higher orders in $\epsilon$. The relations of the master variables in the coupled master equation (27) and those in Eqs. (20)–(22) are given by
@@
$-\frac{8ir_{\rm g}^{4}\omega(75r^{2}+78rr_{\rm g}-10r_{\rm g}^{2})}{15r^{5}}\frac{{\rm d}\tilde{\Psi}^{-}}{{\rm d}r}\Bigg{\}},$ (89)
@@ Appendix B Perturbative eigenvalues of symmetric matrix
In this section, we study the eigenvalues of symmetric matrices that are diagonal at the lowest order but have non-diagonal components perturbatively. Since the characteristics of the perturbative spectrum of eigenvalues of symmetric matrices are similar to those of the spectrum of quasinormal frequencies of black holes, this study is useful for understanding the results of Sec. IV.
@@ B.2 $2\times 2$ case: degenerate
Substituting $\lambda=2+\Delta\lambda$ into Eq. (95), we obtain $(\Delta\lambda)^{2}-\delta^{2}=0$. This implies that the eigenvalues are $\lambda=2\pm\delta$. For symmetric matrices that are degenerate at the lowest order, the correction of the eigenvalue appears at ${\cal O}(\delta)$. This is because the leading-correction term of $\Delta\lambda$ in the characteristic equation is ${\cal O}((\Delta\lambda)^{2})$ if the lowest-order spectra are degenerate.
@@ B.3 $3\times 3$ case
If $p$ takes different values when $b_{1}=0$ and when $b_{2}=0$ while keeping the condition $b_{1}b_{2}+b_{3}=0$, the branches of Eqs. (102) and (103) exchange for $b_{1}=0$ and $b_{2}=0$ cases. This phenomenon can be seen in the QNM spectra for our setup.
@@ Appendix C Static solutions as tidal response for $\ell=2$
In Sec. VI, we discussed the tidal response from the new even-odd couplings. In this section, we present explicit forms of the static solutions with the regularity condition at the horizon for $\ell=2,\omega=0$ in the master equations (27). To obtain perturbative solutions in $\epsilon$, we expand the solutions around the lowest order solutions up to ${\cal O}(\epsilon^{2})$ as Eq. (69). The lowest-order solutions $\tilde{\Psi}^{\rm s}_{\rm Sch},~{}\tilde{\Psi}^{-}_{\rm Sch}$, and $\tilde{\Psi}^{+}_{\rm Sch}$ are given by Eqs. (66)–(68), which corresponds to the Schwarzschild case. We derive the correction terms by solving Eq. (27) with the ansatz in Eq. (69).
C.1 Scalar field-led case
@@
Through the scalar-odd coupling $V^{\rm s-}_{\rm 1st}$ and scalar-even coupling $V^{\rm s+}_{\rm 1st}$ and the self-scalar coupling $V^{\rm s}_{\rm 2nd}$ in Eq. (31), the higher-order solutions are given by
@@ C.2 Odd-led case
Let us consider solutions led by the odd parity external tidal field. The lowest-order growing odd solution is given by Eq. (67)
@@ C.3 Even-led case
We study solutions led by the even parity external tidal field. The lowest-order growing solution is given by Eq. (68) as